Ferdie was having a big party and he was excited. He had been practicing his opening lines all week and as soon as he he got to the part, 20 girls at the party left. There now remained 2 boys for each girl. A lot of the other boys thought the party was pathetic, so 20 of the boys also left. There were now 3 times as many girls as boys. How many boys and girls were at the party when Ferdie got there?
Answer by stanbon(57967)

Ferdie was having a big party and he was excited. He had been practicing his opening lines all week and as soon as he he got to the part, 20 girls at the party left. There now remained 2 boys for each girl. A lot of the other boys thought the party was pathetic, so 20 of the boys also left. There were now 3 times as many girls as boys. How many boys and girls were at the party when Ferdie got there?
--------------------------
Let original # of girls be "G"; Let original number of boys be "B".
---------------------
After 20 girls left:
B = 2(G-20)
--------------------------
After 20 boys left:
(G-20) = 3(B-20)
-------------------------
Substitute for "B" and solve for "G":
G-20 = 3(2G-40-20)
G-20 = 3(2G-60)
G-20 = 6G-180
5G = 160
G = 32 (original number of girls)
---
B=2(G-20)
B = 2*12 = 24 (original number of boys)
===========================================
Cheers,
Stan H.